# Minimax Hypothesis Testing for the Bradley-Terry-Luce Model

### v1

@@ V-A Preliminaries
where $\zeta_{1}$ follows by a standard argument and the explanation is provided below. $\zeta_{2}$ is a consequence of Cheeger’s inequality for non-negative matrices in Lemma 5 (see below). $\zeta_{3}$ follows since $\sqrt{1-x}\leq 1-x/2$ for $x\geq 0.$ and since $\phi(R)=\phi((R+R^{\mathrm{T}})/2)=\phi(R^{\mathrm{T}})$ (proof provided below for completeness). Finally, $\zeta_{4}$ follows because $\phi(R)\geq\xi,$ by Assumption 2. Regarding $\zeta_{1}$, observe that since the matrix $R$ is $\frac{1}{2}$-lazy, i.e., $R_{ii}\geq\frac{1}{2}$ for all $i\in[n]$ (as $S_{ii}\geq\frac{1}{2}$), therefore we have
@@
$x^{\mathrm{T}}RR^{\mathrm{T}}x\leq\frac{x^{\mathrm{T}}(R+R^{\mathrm{T}})x}{2}+\|x\|^{2}_{2}\frac{\|2R-I\|_{2}\|2R^{\mathrm{T}}-I\|_{2}}{4}-\frac{\|x\|_{2}^{2}}{4}\stackrel{{\scriptstyle\zeta}}{{\leq}}\frac{x^{\mathrm{T}}(R+R^{\mathrm{T}})}{2}x,$
where $\zeta$ follows since $2R-I$ is a non-negative matrix (as $R$ is $\frac{1}{2}$-lazy) with largest singular value of $1$ and with both left and right singular vectors $\sqrt{\pi}$, which gives $\|2R-I\|_{2}\leq 1$. This gives
@@
${\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}D_{u}RD_{v}{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\Pi^{1/2}(\Pi^{1/2}S\Pi^{-1/2})\Pi^{1/2}{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\Pi S{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\Pi S({\mathbf{1}}_{n}-{\mathbf{1}}_{\mathcal{S}})={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\pi-{\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\Pi S{\mathbf{1}}_{\mathcal{S}},$
@@
${\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}D_{u}R^{\mathrm{T}}D_{v}{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}$ $={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\Pi^{1/2}(\Pi^{-1/2}S^{\mathrm{T}}\Pi^{1/2})\Pi^{1/2}{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}S^{\mathrm{T}}\Pi{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}S^{\mathrm{T}}\Pi({\mathbf{1}}_{n}-{\mathbf{1}}_{\mathcal{S}})$
$={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\pi-{\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}S^{\mathrm{T}}\Pi{\mathbf{1}}_{\mathcal{S}}={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\pi-{\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\Pi S{\mathbf{1}}_{\mathcal{S}}={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}D_{u}RD_{v}{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}.$
@@
Lemma 5 (Cheeger Inequalities for Non-negative Matrices Satisfying Detailed Balance [63, Theorem 15]).
Consider a non-negative matrix $M$ with a Perron-Frobenius eigenvalue of $1$ and positive left and right eigenvectors $u$ and $v$. Assume that $M$ satisfies the condition of detailed balance, i.e., $D_{u}MD_{v}=D_{v}M^{\mathrm{T}}D_{u}$ where $D_{u}=\operatorname{{diag}}(u)$ and $D_{v}=\operatorname{{diag}}(v)$. Then, the following inequalities hold:
@@
$1-\|\Pi^{1/2}S\Pi^{-1/2}-\sqrt{\pi}\sqrt{\pi}^{\mathrm{T}}\|_{2}-\|{\hat{S}}-S\|_{\pi}$ $\geq\frac{1}{4}\xi^{2}-5\sqrt{\frac{h\log n}{kd_{\max}}}\geq\frac{\xi^{2}}{8},$
@@
Now, using Corollary 1 to bound the denominator term in Lemma 1 and following the same procedure in [11, Theorem 9], we obtain the following $\ell^{2}$-error bound.
@@
Under the pairwise comparison model discussed in Section II such that Assumptions 2 and 4 holds and for $k\geq\max\big{\{}1,\frac{ch\log n}{d_{\max}\xi^{4}}\big{\}}$ and $d_{\max}\geq\log n$, the following bound holds:
$\|\pi-{\hat{\pi}}\|_{2}\leq\frac{C}{\xi^{2}}\sqrt{\frac{h}{kd_{\max}}}\sqrt{n}\|\pi\|_{\infty}\leq\frac{c_{2}\|\pi\|_{2}}{\sqrt{kd_{\max}}},$
with probability at least $1-O(n^{-3})$ for some constants $c,C,c_{2}$ independent of $n,d_{\max},k$ (but dependent on $h$ and $\xi$).
Now utilizing the results developed above the rest of the proof follows using similar arguments in [11, Theorem 9] and is provided in Appendix C for completeness.
@@ V-B Proof of Theorem 2
${\mathbb{E}}[\hat{p}_{ij}]=p_{ij}\ \text{ and }\ {\mathrm{var}}(\hat{p}_{ij})=\frac{p_{ij}(1-p_{ij})}{k_{ij}},$ (52)
${\mathbb{E}}[\hat{Y}_{ij}]=p_{ij}^{2}\ \text{ and }{\mathrm{var}}(\hat{Y}_{ij})=\frac{-2\left(2k_{ij}-3\right)p_{ij}^{4}+4(k_{ij}-2)p_{ij}^{3}+2p_{ij}^{2}}{k_{ij}(k_{ij}-1)}.$
@@
$T_{1}$ $\triangleq\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)^{2}-\left(\pi_{i}+\pi_{j}\right)^{2}\right)\left(\hat{Y}_{ij}-p_{ij}^{2}\right)-2\left(\hat{\pi}_{j}\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)-\pi_{j}\left(\pi_{i}+\pi_{j}\right)\right)\left(\hat{p}_{ij}-p_{ij}\right),$ (53)
$T_{2}$ $\triangleq\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)^{2}-\left(\pi_{i}+\pi_{j}\right)^{2}\right)p_{ij}^{2}+\hat{\pi}_{j}^{2}-\pi_{j}^{2}-2\left(\hat{\pi}_{j}\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)-\pi_{j}\left(\pi_{i}+\pi_{j}\right)\right)p_{ij},$ (54)
@@
The following bounds hold on $T_{1},T_{2}$.
  • There exist constants $c_{0},\tilde{c}_{0},\hat{c}_{0}$ such that the following tail bound holds for $|T_{1}|$
    $\mathbb{P}\left(|T_{1}|\geq c_{0}\frac{n\|\pi\|_{\infty}^{2}}{k}+\frac{\tilde{c}_{0}n\|\pi\|_{\infty}^{2}}{k}\left(\left(\frac{(\log n)^{3}}{d_{\max}}\right)^{1/4}+\sqrt{\frac{t}{d_{\max}}}+\left(\frac{t^{4}}{kd_{\max}}\right)^{1/6}+\sqrt{\frac{t^{2}}{kd_{\max}}}\right)\right)\leq 16ne^{-t}+{\frac{\hat{c}_{0}}{n^{3}}}.$ (56)
@@
    $|T_{1}|\leq\frac{c_{\alpha}n\|\pi\|_{\infty}^{2}}{k}.$ (57)
@@
    $|T_{2}|\leq\frac{c_{\beta}\sqrt{n}\|\pi\|_{\infty}}{\sqrt{k}}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}+c_{\gamma}\frac{n\|\pi\|^{2}_{\infty}}{k}.$ (58)
  • If $d_{\max}\geq\log n$, the constant $c_{\alpha}$ and $c_{\gamma}$ defined above scale as $O(\sqrt{\log n})$
@@
The following bounds hold for the mean and variance of $T_{3}$ as defined in Eq. 55.
@@
    ${\mathrm{var}}(T_{3})\leq\frac{4\|\pi\|^{2}_{\infty}}{k}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}^{2}+\frac{4nd_{\max}}{k^{2}}\|\pi\|_{\infty}^{4}.$
@@
$\{T\geq t\}$ =($=\left(\{T\geq t\}\cap\{T\leq T^{\prime}\}\right)\cup\left(\{T\geq t\}\cap\{T>T^{\prime}\}\right)$
=($=\left(\{t\leq T\leq T^{\prime}\}\right)\cup\left(\{T\geq t\}\cap\{T>T^{\prime}\}\right).$
Let $T^{\prime}=(c_{\alpha}+c_{\gamma})\frac{\|\pi\|^{2}_{\infty}n}{k}+T_{3}.$ Then, we have
${\mathbb{P}}_{H_{0}}(T\geq t)$ $\leq{\mathbb{P}}_{H_{0}}(T^{\prime}\geq t)+{\mathbb{P}}_{H_{0}}(T>T^{\prime})$ (59)
$\leq{\mathbb{P}}_{H_{0}}\left(T_{3}\geq t-(c_{\alpha}+c_{\gamma})n\|\pi\|^{2}_{\infty}/{k}\right)+O\bigg{(}\frac{1}{n^{3}}\bigg{)}$
$\stackrel{{\scriptstyle\zeta}}{{\leq}}\frac{{\mathrm{var}}_{H_{0}}(T_{3})}{{\mathrm{var}}_{H_{0}}(T_{3})+(t-(c_{\alpha}+c_{\gamma})n\|\pi\|^{2}_{\infty}/{k})^{2}}+O\bigg{(}\frac{1}{n^{3}}\bigg{)}$
$\leq\frac{4nd_{\max}\|\pi\|^{4}_{\infty}/k}{4nd_{\max}\|\pi\|^{4}_{\infty}/(k)+16nd_{\max}\|\pi\|^{4}_{\infty}/k}+O\bigg{(}\frac{1}{n^{3}}\bigg{)}\leq\frac{1}{4},$
where $\zeta$ follows from one-sided Chebyshev’s inequality [80], and in the last inequality we have substituted
$t=4\sqrt{nd_{\max}}\|\pi\|^{2}_{\infty}/{k}+(c_{\alpha}+c_{\gamma})\frac{n\|\pi\|^{2}_{\infty}}{k}.$ (60)
Similarly, under hypothesis $H_{1}$, for some random variable $T^{\prime}$, we have
$\{T\leq t\}$ =($=\left(\{T\leq t\}\cap\{T\leq T^{\prime}\}\right)\cup\left(\{T\leq t\}\cap\{T\geq T^{\prime}\}\right)$
=($=\left(\{T\leq t\}\cap\{T\leq T^{\prime}\}\right)\cup\left(\{T^{\prime}\leq T\leq t\}\right).$
@@
$\Delta_{T}=\frac{c_{\beta}\sqrt{n}\|\pi\|_{\infty}}{\sqrt{k}}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}+(c_{\alpha}+c_{\gamma})\frac{n\|\pi\|^{2}_{\infty}}{k}.$
Therefore, we can bound ${\mathbb{P}}_{H_{1}}(T\leq t)$ as
${\mathbb{P}}_{H_{1}}(T\leq t)$ $\leq{\mathbb{P}}_{H_{1}}(T^{\prime}\leq t)+{\mathbb{P}}_{H_{1}}(T^{\prime}\leq T)\leq{\mathbb{P}}_{H_{1}}\left(T_{3}\leq t+\Delta_{t}\right)+O\bigg{(}\frac{1}{n^{3}}\bigg{)}$ (61)
$\leq\frac{{\mathrm{var}}_{H_{1}}(T_{3})}{{\mathrm{var}}_{H_{1}}(T_{3})+({\mathbb{E}}_{H_{1}}[T_{3}]-t-\Delta_{t})^{2}}+O\bigg{(}\frac{1}{n^{3}}\bigg{)}\stackrel{{\scriptstyle\zeta}}{{\leq}}1/4,$
@@
$4{\mathrm{var}}_{H_{1}}(T_{3})\leq({\mathbb{E}}_{H_{1}}[T_{3}]-t-\Delta_{t})^{2}.$
@@
$2\left(\frac{2\|\pi\|_{\infty}}{\sqrt{k}}D+\frac{2\sqrt{nd_{\max}}\|\pi\|^{2}_{\infty}}{{k}}\right)\leq\bigg{(}D^{2}-\frac{c_{\beta}D\sqrt{n}\|\pi\|_{\infty}}{\sqrt{k}}-\frac{4\sqrt{nd_{\max}}\|\pi\|^{2}_{\infty}}{{k}}-2(c_{\alpha}+c_{\gamma})\frac{n\|\pi\|^{2}_{\infty}}{k}\bigg{)}.$
Substituting $D=\epsilon n\|\pi\|_{\infty}$, and cancelling out $\|\pi\|_{\infty}^{2}$ we obtain the following equivalent condition
$\frac{4n\epsilon}{\sqrt{k}}+\frac{4n}{k}\sqrt{\frac{d_{\max}}{n}}\leq\epsilon^{2}n^{2}-\frac{c_{\beta}\epsilon n^{1.5}}{\sqrt{k}}-\frac{4n}{k}\sqrt{\frac{d_{\max}}{n}}-2(c_{\alpha}+c_{\gamma})\frac{n}{k}.$ (62)
Next, substituting $\epsilon=a/\sqrt{nk}$, we get
$\frac{4a}{\sqrt{n}}+4\sqrt{\frac{d_{\max}}{n}}\leq a^{2}-c_{\beta}a-4\sqrt{\frac{d_{\max}}{n}}-4(c_{\alpha}+c_{\gamma}).$ (63)
Clearly, the condition in (63) is satisfied for sufficiently a large constant $a_{0}$ which gives
$nk\epsilon^{2}\geq a_{0}\implies\epsilon\geq\frac{a_{0}}{\sqrt{nk}}.$
[…]
$\varepsilon_{\mathsf{c}}^{2}\leq O\left(\frac{1}{nk}\right).$

### v2

@@ V-A Preliminaries
where $\zeta_{1}$ follows by a standard argument and the explanation is provided below. $\zeta_{2}$ is a consequence of the Cheeger inequality for non-negative matrices in Lemma 5 (see below). $\zeta_{3}$ follows since $\sqrt{1-x}\leq 1-x/2$ for $x\geq 0$, and since $\phi(R)=\phi((R+R^{\mathrm{T}})/2)=\phi(R^{\mathrm{T}})$ (the proof is provided below for completeness). Finally, $\zeta_{4}$ follows because $\phi(R)=\phi(S)\geq\xi,$ by Assumption 1. Regarding $\zeta_{1}$, observe that since the matrix $R$ is $\frac{1}{2}$-lazy, i.e., $R_{ii}\geq\frac{1}{2}$ for all $i\in[n]$ (as $S_{ii}\geq\frac{1}{2}$), therefore we have
@@
$x^{\mathrm{T}}RR^{\mathrm{T}}x$ $\leq\frac{x^{\mathrm{T}}(R+R^{\mathrm{T}})x}{2}$
$\ \ \ \ \ \ \ \ +\|x\|^{2}_{2}\frac{\|(2R-I)(2R^{\mathrm{T}}-I)\|_{2}}{4}-\frac{\|x\|_{2}^{2}}{4}$
$\stackrel{{\scriptstyle\zeta}}{{\leq}}\frac{x^{\mathrm{T}}(R+R^{\mathrm{T}})x}{2},$
where $\zeta$ follows since $(2R-I)(2R^{\mathrm{T}}-I)$ is a doubly non-negative matrix (as $R$ is $\frac{1}{2}$-lazy) with largest eigenvalue (and, hence, singular value) of $1$ with eigenvector $\sqrt{\pi}$, which gives $\|(2R-I)(2R^{\mathrm{T}}-I)\|_{2}=1$. This gives
@@
${\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}D_{u}RD_{v}{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}$ $={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\Pi^{1/2}(\Pi^{1/2}S\Pi^{-1/2})\Pi^{1/2}{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}$
$={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\Pi S{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\Pi S({\mathbf{1}}_{n}-{\mathbf{1}}_{\mathcal{S}})$
$={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\pi-{\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\Pi S{\mathbf{1}}_{\mathcal{S}},$
@@
${\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}D_{u}R^{\mathrm{T}}D_{v}{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}$ $={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\Pi^{1/2}(\Pi^{-1/2}S^{\mathrm{T}}\Pi^{1/2})\Pi^{1/2}{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}$
$={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}S^{\mathrm{T}}\Pi{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}S^{\mathrm{T}}\Pi({\mathbf{1}}_{n}-{\mathbf{1}}_{\mathcal{S}})$
$={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\pi-{\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}S^{\mathrm{T}}\Pi{\mathbf{1}}_{\mathcal{S}}$
$={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\pi-{\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}\Pi S{\mathbf{1}}_{\mathcal{S}}={\mathbf{1}}_{\mathcal{S}}^{\mathrm{T}}D_{u}RD_{v}{\mathbf{1}}_{{\mathcal{S}}^{\mathrm{c}}}.$
@@
Lemma 5 (Cheeger Inequalities for Non-negative Matrices Satisfying Detailed Balance [MehtaSchulman, Theorem 15]).
Consider a non-negative matrix $M$ with a Perron-Frobenius eigenvalue of 11 and positive left and right eigenvectors $u$ and $v$. Assume that $M$ satisfies the condition of detailed balance, i.e., $D_{u}MD_{v}=D_{v}M^{\mathrm{T}}D_{u}$ where $D_{u}=\operatorname{{diag}}(u)$ and $D_{v}=\operatorname{{diag}}(v)$. Then, the following inequalities hold:
@@
$1-\|\Pi^{1/2}S\Pi^{-1/2}$- $\sqrt{\pi}\sqrt{\pi}^{\mathrm{T}}\|_{2}-\|{\hat{S}}-S\|_{\pi}$
$\geq\frac{1}{4}\xi^{2}-5\sqrt{\frac{h\log n}{kd_{\max}}}\geq\frac{\xi^{2}}{8},$
@@
Now, using Corollary 1 to bound the denominator term in Lemma 1 and following the same procedure in [Chenetal2019, Theorem 9], we obtain the following $\ell^{2}$-error bound.
@@
Under the pairwise comparison model discussed in Section II such that Assumptions 1 and 2 holds and for $k\geq\max\big\{1,\frac{ch\log n}{d_{\max}\xi^{4}}\big\}$ and $d_{\max}\geq\log n$, the following bound holds:
$\|\pi-{\hat{\pi}}\|_{2}\leq\frac{C}{\xi^{2}}\sqrt{\frac{h}{kd_{\max}}}\sqrt{n}\|\pi\|_{\infty}$
with probability at least $1-O(n^{-3})$ for some constants $c,C$ independent of $n,d_{\max},k$.
Utilizing the results developed above, the rest of the proof follows using similar arguments in [Chenetal2019, Theorem 9] and is provided in Appendix C for completeness.
@@ V-B Proof of Theorem 2
${\mathbb{E}}[\hat{p}_{ij}]=p_{ij}\ \text{ and }\ {\mathrm{var}}(\hat{p}_{ij})=\frac{p_{ij}(1-p_{ij})}{k_{ij}},$
${\mathbb{E}}[\hat{Y}_{ij}]=p_{ij}^{2}\ \text{ and }$
${\mathrm{var}}(\hat{Y}_{ij})=\frac{-2\left(2k_{ij}-3\right)p_{ij}^{4}+4(k_{ij}-2)p_{ij}^{3}+2p_{ij}^{2}}{k_{ij}(k_{ij}-1)}.$ (50)
@@
$T_{1}$ $\triangleq\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)^{2}-\left(\pi_{i}+\pi_{j}\right)^{2}\right)\left(\hat{Y}_{ij}-p_{ij}^{2}\right)$
$\ \ \ \ \ \ \ \ \ -2\left(\hat{\pi}_{j}\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)-\pi_{j}\left(\pi_{i}+\pi_{j}\right)\right)\left(\hat{p}_{ij}-p_{ij}\right),$ (51)
$T_{2}$ $\triangleq\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{E}}:\\ i\neq j\end{subarray}}\left(\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)^{2}-\left(\pi_{i}+\pi_{j}\right)^{2}\right)p_{ij}^{2}+\hat{\pi}_{j}^{2}-\pi_{j}^{2}$
$\ \ \ \ \ \ \ \ \ -2\left(\hat{\pi}_{j}\left(\hat{\pi}_{i}+\hat{\pi}_{j}\right)-\pi_{j}\left(\pi_{i}+\pi_{j}\right)\right)p_{ij},$ (52)
@@
Under the assumptions of Theorem 2, the following bounds hold on $T_{1},T_{2}$:
  1. 1.
    There exist constants $c_{0},\tilde{c}_{0},\hat{c}_{0}$ such that the following tail bound holds for $|T_{1}|$:
[…]
    $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \leq 16ne^{-t}+{\frac{\hat{c}_{0}}{n^{3}}}.$ (54)
  2. 2.
@@
    $|T_{1}|\leq\frac{c_{\alpha}n\|\pi\|_{\infty}^{2}\sqrt{h}}{k\xi^{2}}.$ (55)
  3. 3.
@@
    $|T_{2}|$ $\leq\frac{c_{\beta}\sqrt{n}\|\pi\|_{\infty}\sqrt{h}}{\sqrt{k}\xi^{2}}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}$
    $\quad\,+c_{\gamma}\frac{n\|\pi\|^{2}_{\infty}h}{k\xi^{4}}.$ (56)
  4. 4.
    If $d_{\max}\geq\log n$, the constants $c_{\alpha}$ and $c_{\gamma}$ defined above scale as $O(\sqrt{\log n})$.
@@
The following bounds hold for the mean and variance of $T_{3}$ as defined in Eq. 53:
@@
    ${\mathrm{var}}(T_{3})$ $\leq\frac{4\|\pi\|^{2}_{\infty}}{k}\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}^{2}$
    $\quad\,+\frac{4nd_{\max}}{k^{2}}\|\pi\|_{\infty}^{4}.$
@@
$\{T\geq t\}$ $=\left(\{T\geq t\}\cap\{T\leq T^{\prime}\}\right)\cup\left(\{T\geq t\}\cap\{T>T^{\prime}\}\right)$
$=\left(\{t\leq T\leq T^{\prime}\}\right)\cup\left(\{T\geq t\}\cap\{T>T^{\prime}\}\right).$
Define $\tilde{c}_{\alpha}=c_{\alpha}\sqrt{h}/\xi^{2}$, $\tilde{c}_{\beta}=c_{\beta}\sqrt{h}/\xi^{2}$ and $\tilde{c}_{\gamma}=c_{\gamma}h/\xi^{4}$. Let $T^{\prime}=(\tilde{c}_{\alpha}+\tilde{c}_{\gamma})\frac{\|\pi\|^{2}_{\infty}n}{k}+T_{3}.$ Then, we have (cf. [RastogiBalakrishnanShahAarti2022])
${\mathbb{P}}_{H_{0}}(T\geq t)\leq{\mathbb{P}}_{H_{0}}(T^{\prime}\geq t)+{\mathbb{P}}_{H_{0}}(T>T^{\prime})$
$\leq{\mathbb{P}}_{H_{0}}\left(T_{3}\geq t-(\tilde{c}_{\alpha}+\tilde{c}_{\gamma})n\|\pi\|^{2}_{\infty}/{k}\right)+O\bigg(\frac{1}{n^{3}}\bigg)$
$\stackrel{{\scriptstyle\zeta_{1}}}{{\leq}}\frac{{\mathrm{var}}_{H_{0}}(T_{3})}{{\mathrm{var}}_{H_{0}}(T_{3})+(t-(\tilde{c}_{\alpha}+\tilde{c}_{\gamma})n\|\pi\|^{2}_{\infty}/{k})^{2}}+O\bigg(\frac{1}{n^{3}}\bigg)$
$\stackrel{{\scriptstyle\zeta_{2}}}{{\leq}}\frac{4nd_{\max}\|\pi\|^{4}_{\infty}/k^{2}}{4nd_{\max}\|\pi\|^{4}_{\infty}/k^{2}+16nd_{\max}\|\pi\|^{4}_{\infty}/k^{2}}+O\bigg(\frac{1}{n^{3}}\bigg)$
$\stackrel{{\scriptstyle\zeta_{3}}}{{\leq}}\frac{1}{4},$
where $\zeta_{1}$ follows from one-sided Chebyshev’s inequality [HDP], $\zeta_{3}$ follows for $n$ large enough, and in $\zeta_{2}$ we have substituted
$t=4\frac{\sqrt{nd_{\max}}\|\pi\|^{2}_{\infty}}{k}+(\tilde{c}_{\alpha}+\tilde{c}_{\gamma})\frac{n\|\pi\|^{2}_{\infty}}{k},$
which is clearly upper bounded by the theoretical threshold $\mathsf{thr}$} (independent of $\pi$) of our test given by
${\mathsf{thr}}$ $\triangleq 4\frac{\sqrt{d_{\max}}h^{2}}{n^{3/2}k}+(\tilde{c}_{\alpha}+\tilde{c}_{\gamma})\frac{h^{2}}{nk}$
$=4\frac{\sqrt{d_{\max}}h^{2}}{n^{3/2}k}+\left(\frac{{c}_{\alpha}\sqrt{h}}{\xi^{2}}+\frac{{c}_{\gamma}h}{\xi^{4}}\right)\frac{h^{2}}{nk}.$ (57)
Similarly, under hypothesis $H_{1}$, for some random variable $T^{\prime}$ and $t$, we have
$\{T<t\}$ $=\left(\{T<t\}\cap\{T<T^{\prime}\}\right)\cup\left(\{T<t\}\cap\{T\geq T^{\prime}\}\right)$
$=\left(\{T<t\}\cap\{T<T^{\prime}\}\right)\cup\left(\{T^{\prime}\leq T<t\}\right).$
@@
$\Delta_{T}=\frac{\tilde{c}_{\beta}\sqrt{n}\|\pi\|_{\infty}}{\sqrt{k}}$ $\|\Pi P+P\Pi-{\mathcal{P}_{\mathcal{E}}}(\mathbf{1}_{n}\pi^{\mathrm{T}})\|_{\mathrm{F}}$
$+(\tilde{c}_{\alpha}+\tilde{c}_{\gamma})\frac{n\|\pi\|^{2}_{\infty}}{k}.$
Therefore, we can bound ${\mathbb{P}}_{H_{1}}(T<{\mathsf{thr}})$ as
${\mathbb{P}}_{H_{1}}(T<{\mathsf{thr}})\leq{\mathbb{P}}_{H_{1}}(T^{\prime}<{\mathsf{thr}})+{\mathbb{P}}_{H_{1}}(T^{\prime}>T)$
$\leq{\mathbb{P}}_{H_{1}}\left(T_{3}<{\mathsf{thr}}+\Delta_{T}\right)+O\bigg(\frac{1}{n^{3}}\bigg)$
$\leq\frac{{\mathrm{var}}_{H_{1}}(T_{3})}{{\mathrm{var}}_{H_{1}}(T_{3})+({\mathbb{E}}_{H_{1}}[T_{3}]-{\mathsf{thr}}-\Delta_{T})^{2}}+O\bigg(\frac{1}{n^{3}}\bigg)\stackrel{{\scriptstyle\zeta}}{{\leq}}\frac{1}{4},$
@@
$4{\mathrm{var}}_{H_{1}}(T_{3})\leq({\mathbb{E}}_{H_{1}}[T_{3}]-{\mathsf{thr}}-\Delta_{T})^{2}.$
@@
$2\left(\frac{2h}{n\sqrt{k}}D+\frac{2\sqrt{nd_{\max}}h^{2}}{n^{2}{k}}\right)\leq$
$\ \ \ \ \ \ \ \ \ \ \ \ \bigg(D^{2}-\frac{\tilde{c}_{\beta}Dh}{\sqrt{nk}}-\frac{4h^{2}\sqrt{d_{\max}}}{n^{3/2}{k}}-2(\tilde{c}_{\alpha}+\tilde{c}_{\gamma})\frac{h^{2}}{nk}\bigg).$
Substituting $D=\epsilon n\|\pi\|_{\infty}=a/\sqrt{nk}$, we obtain the following equivalent condition:
$\frac{4ah}{n^{3/2}k}+\frac{4h^{2}\sqrt{d_{\max}}}{n^{3/2}k}$ $\leq\frac{a^{2}n\|\pi\|_{\infty}^{2}}{k}-\frac{\tilde{c}_{\beta}ah}{nk}$
$-4\frac{\sqrt{d_{\max}}h^{2}}{n^{3/2}k}-2(\tilde{c}_{\alpha}+\tilde{c}_{\gamma})\frac{h^{2}}{nk}.$ (58)
Using that $n\|\pi\|_{\infty}\leq 1$, This is equivalent to the following condition:
$\frac{4a}{hn^{1/2}}+4\sqrt{\frac{d_{\max}}{n}}\leq\frac{a^{2}}{h^{2}}-\tilde{c}_{\beta}\frac{a}{h}-4\sqrt{\frac{d_{\max}}{n}}-2(\tilde{c}_{\alpha}+\tilde{c}_{\gamma}).$ (59)
Now, we can rewrite the above equation as a polynomial in $a$ as:
$\frac{a^{2}}{h^{2}}$- $\left(\frac{\sqrt{h}}{\xi^{2}}c_{\beta}+\frac{4}{\sqrt{n}}\right)\frac{a}{h}$
$\ \ \ \ \ \ \ \ -\left(8\sqrt{\frac{d_{\max}}{n}}+2\left({c}_{\alpha}\frac{\sqrt{h}}{\xi^{2}}+{c}_{\gamma}\frac{h}{\xi^{4}}\right)\right)\geq 0.$ (60)
Utilizing the fact that $\xi\leq 1$ and $h\geq 1$, the condition in (60) is clearly satisfied for a sufficiently large constant $a_{0}$ such that $a\geq a_{0}h^{3/2}/\xi^{2}$, which gives
$n\|\pi\|_{\infty}\epsilon\geq\frac{a_{0}h^{3/2}}{\xi^{2}\sqrt{nk}}\implies\epsilon\geq\frac{a_{0}\sqrt{h^{3}}}{\sqrt{nk\xi^{4}}}.$
Thus, for sufficiently large constant $a_{0}$, we have demonstrated that the sum of type I and type II errors is bounded by $\frac{1}{2}$. Combining this result with the definition of the critical threshold in Eq. 21, we obtain the following bound on $\varepsilon_{\mathsf{c}}$:
$\varepsilon_{\mathsf{c}}^{2}\leq O\left(\frac{h^{3}}{nk\xi^{4}}\right).$